OpenStudy (anonymous):

9/11 X 5/6 i get 45/66 i cant simplify any much more but still not the answer

hero (hero):

\[\frac{9}{11} \times \frac{5}{6}\]

hero (hero):

You can cross simplify the 9 and the 6

hero (hero):

If you do that, you'll get: \[\frac{3}{11} \times \frac{5}{2}\] Now try multiplying that

OpenStudy (anonymous):

15/22

OpenStudy (calculusxy):

The GCF of 45 and 66 is 3. So divide 45 by 3 and 66 by 3. When doing so, you will get 45/3=15 and 66/3=22. So the final answer should be 15/22.

OpenStudy (calculusxy):

So your right!

hero (hero):

Because, you know, \[\frac{9}{6} = \frac{3}{2}\]

OpenStudy (anonymous):

ahhh thats the trick thank you

hero (hero):

Its best to try to reduce the initial expressions as much as possible BEFORE multiplying. You don't want to end up multiplying and getting bigger numbers than necessary.

OpenStudy (anonymous):

il note that down thanks hero